# Descend or Rewind? Stochastic Gradient Descent Unlearning

### v1

@@ 4 Analyses
We first discuss general assumptions required for this work. We assume that the loss function $\ell$ is differentiable and Lipschitz smooth (Assumption 1), a standard requirement for the analysis of SGD algorithms.
Assumption 1.
@@
In addition to nonconvex functions, we analyze the cases of convex and strongly convex functions, defined below in Assumptions 2 and 3.
Assumption 2.
@@
Assumption 3.
For all $z\in\mathcal{Z}$, the function $\ell(z;\theta)$ is strongly convex with parameter $\mu$ such that for any $\theta_{1},\theta_{2}\in\mathbb{R}^{d}$,
@@
Now we broadly describe our proof approach. Our analysis relies on carefully tracking the (P)SGD trajectories on $\mathcal{L}_{\mathcal{D}}$ and $\mathcal{L}_{\mathcal{D}^{\prime}}$ and bounding the outputs of training on $\mathcal{L}_{\mathcal{D}^{\prime}}$ and unlearning on $\mathcal{L}_{\mathcal{D}}$. Specifically, we have that
  • $\{\theta_{t}\}_{t=0}^{T}$ represents the learning iterates on $\mathcal{L}_{\mathcal{D}}$, starting from $\theta_{0}$,
  • $\{\theta^{\prime}_{t}\}_{t=0}^{T}$ represents the learning iterates on $\mathcal{L}_{\mathcal{D^{\prime}}}$, where $\theta^{\prime}_{0}=\theta_{0}$,
  • $\{\theta^{\prime\prime}_{t}\}_{t=0}^{K}$ represents the unlearning iterates on $\mathcal{L}_{\mathcal{D}^{\prime}}$, where for D2D we have $\theta^{\prime\prime}_{0}=\theta_{T}$ and for R2D we have at $\theta^{\prime\prime}_{0}=\theta_{T-K}$.
The goal of our work is to bound the distance between $\theta^{\prime}_{T}$ and $\theta^{\prime\prime}_{K}$. As highlighted in the introduction, our analysis relies on a key coupling argument, which enables the application of a relaxed Gaussian mechanism. Basically, we observe that the definition of $(\epsilon,\delta)$-unlearning (Definition 2) involves only the marginal distributions of $\theta^{\prime}_{T}$ and $\theta^{\prime\prime}_{K}$, we are free, for the purpose of analysis, to consider any coupling of these marginals that is convenient; in particular, we choose a coupling that minimizes the expected distance between them. Like in DP, unlearning involves comparing against a hypothetical output that may not be realized in practice: for DP, the point of comparison is the algorithm output on a dataset with one sample replaced, and for certified unlearning, we compare against the output of retraining from scratch. Therefore, we only need to show that the real output— the output of the algorithm called on the dataset $\mathcal{D}$— is probabilistically indistinguishable from a valid instance of the hypothetical output.
This rationale allows us to couple the randomization (the “coin flips”) of the learning and unlearning algorithms to minimize the distances between $\theta_{t}$ and $\theta^{\prime}_{t}$, as well as $\theta^{\prime}_{t}$ and $\theta^{\prime\prime}_{t}$. Specifically, we can carefully choose the batches sampled at each time step so that they coincide as much as possible. We obtain a sensitivity bound on $\lVert\theta^{\prime}_{T}-\theta^{\prime\prime}_{K}\rVert$ that holds with high probability with respect to the chosen joint distribution; in our work, we combine Markov’s inequality with first or second moment bounds, which are easier to obtain from standard analysis of SGD. We can roll the probability of failure into the $(\epsilon,\delta)$ guarantee of the relaxed Gaussian mechanism, described in the lemma below.
Lemma 1.
(Informal) Let $A$ and $U$ denote randomized algorithms that map datasets to model parameters. For any fixed dataset $\mathcal{D}$, let $x=A(\mathcal{D})$ and $y=U(\mathcal{D})$ denote their random outputs, where their randomness is due to the internal coin flips of the algorithm (e.g., mini-batch sampling). Let $\xi,\xi^{\prime}$ be draws from the Gaussian distribution $\mathcal{N}(0,\sigma^{2})$ where $\sigma\geq\frac{\Sigma\sqrt{2\log(1.25/\delta)}}{\epsilon}$. Suppose $\lVert x-y\rVert\leq\Sigma$ with probability $1-\delta$, where the probability is taken over the internal randomness of $A$ and $U$. Then $x$ and $y$ are $(\epsilon,2\delta)$-indistinguishable.
@@
In particular, the “algorithmic randomness” for SGD refers to the mini-batch indices sampled from a fixed distribution (e.g., uniformly from $\{1,...,n\}$), where this distribution does not depend on the actual data values. Like for the classic Gaussian mechanism, we require that the bound $\Sigma$ also does not encode information about the specific dataset, thereby preventing data leakage. We emphasize that the relaxed Gaussian mechanism differs from $(\epsilon,\delta,\gamma)$-random differential privacy [hallRDP, rubinstein17a], which considers whether $(\epsilon,\delta)$-privacy holds with probability $1-\gamma$ with respect to the sampling of the dataset from a well-behaved data distribution.
@@
Our analysis of the R2D framework relies on characterizing unlearning as biased or disturbed SGD on a gradient system. Specifically, we can characterize the SGD iterates on $\mathcal{L}_{\mathcal{D}}$ as biased SGD iterates on $\mathcal{L}_{\mathcal{D}^{\prime}}$, where the unlearning bias or disturbance at each step is the difference between the gradients on $\mathcal{L}_{\mathcal{D}}$ and $\mathcal{L}_{\mathcal{D}^{\prime}}$. Upon unlearning, we perform unbiased SGD on $\mathcal{L}_{\mathcal{D}^{\prime}}$. Our results on strongly convex, convex, and nonconvex functions align with nonlinear contraction theory [kozachkovgeneralization, sontag2021remarks] that predicts that $\theta_{t}$ and $\theta^{\prime}_{t}$ will stay close for perturbed contracting systems (strongly convex functions), diverge linearly on perturbed semi-contracting systems (convex functions), and diverge exponentially otherwise (nonconvex functions). Like in the original R2D proof [mu2025rewindtodeletecertifiedmachineunlearning], rewinding reverses the accumulation of these disturbances, drawing the unlearning trajectory closer to the retraining trajectory. Moreover, rewinding allows us to completely eliminate the impact of noise during unlearning by coupling. However, because the contraction analysis only holds with respect to a Riemannian metric, we can only achieve a first-moment sensitivity bound for the relaxed Gaussian mechanism.
@@ 4.1 PSGD-R2D: Rewinding with Projected SGD
Theorem 2.
Suppose that the loss function $\ell$ satisfies Assumptions 1, and within the closed, bounded, and convex set $\mathcal{C}\subset\mathbb{R}^{d}$, the gradient of $\ell$ is uniformly bounded by some constant $G\geq 0$ such that for all $z\in\mathcal{Z}$ and $\theta\in\mathcal{C}$,
$\lVert\nabla\ell(z;\theta)\rVert\leq G.$
We implement PSGD-R2D (Algorithms 1 and 2 with projection) with
$\sigma=\frac{\Sigma\sqrt{2\log(1.25/\delta)}}{\epsilon\delta}.$
Then PSGD-R2D is an $(\epsilon,2\delta)$-unlearning algorithm for the following choices of $\Sigma$:
@@
  • If $\ell$ is convex (satisfies Assumption 2), then for $\eta\leq\frac{2}{L}$ we have
@@
  • If $\ell$ is $\mu$-strongly convex (satisfies Assumption 3), then for $\eta\leq\frac{\mu}{L^{2}}$ and $\gamma=\sqrt{1-\eta\mu}$ we have
@@
Our approach relies on coupling the trajectories $\theta_{t}$ and $\theta^{\prime}_{t}$, and uniformly bounding the bias with $G$. Theorem 2 highlights the privacy-utility-complexity tradeoff of PSGD-R2D, through the relationship between $\Sigma$, $\epsilon$ and $K$. We observe that for all settings, $\Sigma$ decreases to zero as $K$ increases from zero to $T$, and it decays exponentially for strongly convex functions and linearly for convex functions. For general nonconvex functions, $\Sigma$ still decreases to zero as $K$ goes to $T$, but at a slower rate. This reflects the contractive, semi-contractive, and expansive properties of gradient algorithms on strongly convex, convex, and nonconvex functions respectively.
@@
If $\ell$ is $\mu$-strongly convex (satisfies Assumption 3), then $\Sigma$ is uniformly upper bounded as follows.
@@
Proof. See Appendix B.5.1.
We note that $\frac{n\mu\Sigma}{2\eta G}+\gamma^{T}\leq 1$, so the equation of $K$ is overall positive. For small $T$, $K$ increases at an approximately linear rate, but as $T$ grows to infinity, $K$ converges to a constant value. This implies that for a given level of of privacy $(\epsilon,\delta)$ and noise, the number of unlearning iterations required is uniformly upper bounded even if the number of training iterations $T$ is arbitrarily large.
@@ 4.2 SGD-R2D: Rewinding with SGD
Now we consider rewinding without projection. A key challenge of analyzing SGD unlearning algorithms on an unbounded domain is that the noise and unlearning bias may also be unbounded. To control the noise, we first require a reasonable assumption that the second moment of the stochastic gradient is relatively bounded, as in Assumption 4.
Assumption 4.
For all datasets $\tilde{\mathcal{D}}\sim\mathcal{Z}$, let $g_{\mathcal{B}}(\theta)$ represent the gradient estimator constructed from a batch $\mathcal{B}\sim\tilde{\mathcal{D}}$ uniformly sampled with replacement. Then the second moment of the gradient estimator is relatively bounded with constants $B,C\geq 0$ as follows:
$\mathbb{E}[\lVert g_{\mathcal{B}}(\theta)\rVert^{2}]\leq\mathbb{E}_{z\sim\tilde{\mathcal{D}}}[\lVert\nabla\ell(z;\theta)\rVert^{2}]\leq B\lVert\nabla\mathcal{L}_{\tilde{\mathcal{D}}}(\theta)\rVert^{2}+C,\qquad\text{for all }\theta\in\mathbb{R}^{d}.$ (7)
For finite-sum problems (1), Assumption 4 is satisfied by strongly convex loss functions and functions that satisfy the Polyak–Łojasiewicz (PL) inequality, as established in [khaled2020bettertheorysgdnonconvex, garrigos2024handbookconvergencetheoremsstochastic, gower2021sgdstructurednonconvexfunctions]. In particular, for batch sampling with replacement, the constant $B$ can be computed from $L$ and the batch size $b$, and the constant $C$ can be estimated from how close the model is to interpolation, where perfect interpolation implies $C=0$.
To handle the unlearning bias, we observe that it can be bounded by the $\chi^{2}$-divergence between the empirical distributions of $\mathcal{D}$ and $\mathcal{D}^{\prime}$, denoted as $P_{\mathcal{D}}$ and $P_{\mathcal{D}^{\prime}}$ respectively. Specifically, we show (Lemma 19) that
$\lVert\nabla\mathcal{L}_{\mathcal{D}^{\prime}}(\theta)-\nabla\mathcal{L}_{\mathcal{D}}(\theta)\rVert^{2}=$ $\lVert\mathbb{E}_{z\sim P_{\mathcal{D}^{\prime}}}[\nabla\ell(z;\theta)]-\mathbb{E}_{z\sim P_{\mathcal{D}}}[\nabla\ell(z;\theta)]\rVert^{2}$
$\leq$ $\chi^{2}(P_{\mathcal{D}^{\prime}}\|P_{\mathcal{D}})\;\cdot\;\mathrm{Var}_{P_{\mathcal{D}}}(\nabla\ell(z;\theta)).$
This convenient bound allows us to exploit existing conditions on the noise like Assumption 4.
@@
Assumption 5 is satisfied if the underlying data distribution $\mathcal{Z}$ is closed and bounded.
@@
Theorem 3.
Suppose that the loss function $\ell$ satisfies Assumptions 1, 4, and 5 and we implement SGD-R2D (Algorithms 1 and 2 without projection) with
$\sigma=\frac{\Sigma\sqrt{2\log(1.25/\delta)}}{\epsilon\delta}.$
Then SGD-R2D is an $(\epsilon,2\delta)$-unlearning algorithm for the following choices of $\Sigma$:
@@
    $\Sigma=\frac{(1+\eta L)^{T}-(1+\eta L)^{K}}{L}\left(3B\left(\frac{2}{\eta}\ell_{\theta_{0}}+L\eta C(T-K)\right)(\frac{3n-m}{n-m})+6(\frac{4n-3m}{n-m})C\right)^{1/2}.$
  • If $\ell$ is convex (satisfies Assumption 2), then for $\eta\leq\frac{2}{L}$ we have
    $\Sigma=\eta(T-K)\left(3B\left(\frac{2}{\eta}\ell_{\theta_{0}}+L\eta C(T-K)\right)(\frac{3n-m}{n-m})+6(\frac{4n-3m}{n-m})C\right)^{1/2}.$
  • If $\ell$ is $\mu$-strongly convex (satisfies Assumption 3), then for $\eta\leq\frac{\mu}{L^{2}}$ and $\gamma=\sqrt{1-\eta\mu}$ we have
    $\Sigma=\frac{\gamma^{K}-\gamma^{T}}{\mu}\left(3B\left(\frac{2}{\eta}\ell_{\theta_{0}}+L\eta C(T-K)\right)(\frac{3n-m}{n-m})+6(\frac{4n-3m}{n-m})C\right)^{1/2}.$
Proof. See Appendix B.6.

### v2

@@ 4 Analyses
We first discuss general assumptions required for this work. We assume that the loss function $\ell$ is differentiable and Lipschitz smooth (Assumption 4.1), standard requirements for the analysis of SGD algorithms.
Assumption 4.1.
@@
In addition to nonconvex functions, we analyze the cases of convex and strongly convex functions, defined below in Assumptions 4.2 and 4.3.
Assumption 4.2.
@@
Assumption 4.3.
For all $z\in\mathcal{Z}$, the function $\ell(z;\theta)$ is $\mu$-strongly convex such that for any $\theta_{1},\theta_{2}\in\mathbb{R}^{d}$,
@@
Now we broadly describe our proof approach. Our analysis relies on carefully tracking the SGD trajectories on $\mathcal{L}_{\mathcal{D}}$ and $\mathcal{L}_{\mathcal{D}^{\prime}}$ and bounding the outputs of training on $\mathcal{L}_{\mathcal{D}^{\prime}}$ and unlearning on $\mathcal{L}_{\mathcal{D}}$. Specifically, we have that $\{\theta_{t}\}_{t=0}^{T}$ and $\{\theta^{\prime}_{t}\}_{t=0}^{T}$ represent the learning iterates on $\mathcal{L}_{\mathcal{D}}$ and $\mathcal{L}_{\mathcal{D^{\prime}}}$ respectively, starting from $\theta^{\prime}_{0}=\theta_{0}$. In addition, $\{\theta^{\prime\prime}_{t}\}_{t=0}^{K}$ represents the unlearning iterates on $\mathcal{L}_{\mathcal{D}^{\prime}}$, where for D2D we have $\theta^{\prime\prime}_{0}=\theta_{T}$ and for R2D we have $\theta^{\prime\prime}_{0}=\theta_{T-K}$. The goal is to bound the distance between $\theta^{\prime}_{T}$ and $\theta^{\prime\prime}_{K}$.
As highlighted in the introduction, our analysis relies on a key coupling argument leading to a sensitivity bound that holds in expectation. We couple the randomization (the “coin flips”) of the learning and unlearning algorithms to minimize the distances between $\theta_{t}$ and $\theta^{\prime}_{t}$, as well as $\theta^{\prime}_{t}$ and $\theta^{\prime\prime}_{t}$, by choosing the batches sampled at each time step so that they coincide as much as possible. This approach is similar to existing proofs in differential privacy (Abadi et al., 2016). However, these works evaluate the deviation on the full datasets $\mathcal{D}$ and $\mathcal{D}^{\prime}$ and combine privacy amplification by subsampling with noise at every step to achieve DP. In contrast, to maintain the output perturbation structure (which is amenable to black-box unlearning), we evaluate the sensitivity over the coupling over all steps, obtaining a sensitivity bound on $\lVert\theta^{\prime}_{T}-\theta^{\prime\prime}_{K}\rVert$ that holds in expectation with respect to the chosen joint distribution. This is combined with Markov’s inequality to yield a tail bound of probability $1-\delta$, producing an $(\varepsilon,2\delta)$ guarantee.
Lemma 4.4.
Let $x$ and $y$ be random variables over some domain $\Omega$, and let $\xi,\xi^{\prime}$ be draws from the Gaussian distribution $\mathcal{N}(0,\sigma^{2})$. Then for $0<\varepsilon\leq 1$ and $\delta>0$, we have the following.
  1. 1.
    If the second moment of the distance between $x$ and $y$ is bounded as $\mathbb{E}[\lVert x-y\rVert^{2}]\leq\Sigma^{2}$, then $X=x+\xi$, $Y=y+\xi^{\prime}$ are $(\varepsilon,2\delta)$-indistinguishable if
    $\sigma=\frac{\Sigma}{\varepsilon}\sqrt{\frac{2\log(1.25/\delta)}{\delta}}.$
  2. 2.
    If the first moment of the distance between $x$ and $y$ is bounded as $\mathbb{E}[\lVert x-y\rVert]\leq\Sigma$, then $X=x+\xi$, $Y=y+\xi^{\prime}$ are $(\varepsilon,2\delta)$-indistinguishable if
    $\sigma=\frac{\Sigma\sqrt{2\log(1.25/\delta)}}{\varepsilon\delta}.$
Proof. See Appendix A.3.
@@
Lemma 4.4 maintains differential privacy because the algorithmic randomness (for SGD, the mini-batch indices sampled uniformly from $\{1,...,n\}$) does not depend on the actual data values. Like classic DP methods, we require that the bound $\Sigma$ also does not encode information about the specific dataset, thereby preventing data leakage. We emphasize that our approach differs from $(\varepsilon,\delta,\gamma)$-random differential privacy (Hall et al., 2013; Rubinstein and Aldà, 2017), which considers whether $(\varepsilon,\delta)$-privacy holds with probability $1-\gamma$ with respect to the sampling of the dataset from a well-behaved data distribution.
@@
Our analysis of the R2D framework relies on characterizing unlearning as biased or disturbed SGD on a gradient system. The SGD iterates on $\mathcal{L}_{\mathcal{D}}$ are biased SGD iterates on $\mathcal{L}_{\mathcal{D}^{\prime}}$, where the unlearning bias is the difference between the gradients on $\mathcal{L}_{\mathcal{D}}$ and $\mathcal{L}_{\mathcal{D}^{\prime}}$. Upon unlearning, we perform unbiased SGD on $\mathcal{L}_{\mathcal{D}^{\prime}}$. Our results on strongly convex, convex, and nonconvex functions align with nonlinear contraction theory (Kozachkov et al., 2022; Sontag, 2021) that predicts that $\theta_{t}$ and $\theta^{\prime}_{t}$ will stay close for perturbed contracting systems (strongly convex functions), diverge linearly on perturbed semi-contracting systems (convex functions), and diverge exponentially otherwise (nonconvex functions). Like in (Mu and Klabjan, 2025), rewinding reverses the accumulation of these disturbances, drawing the unlearning trajectory closer to the retraining trajectory. Moreover, rewinding allows us to completely eliminate the impact of noise during unlearning by coupling. However, because the contraction analysis only holds with respect to a Riemannian metric, we can only achieve a first-moment sensitivity bound.
@@ 4.1 PSGD-R2D: Rewinding with Projected SGD
Theorem 4.5.
Suppose that the loss function $\ell$ satisfies Assumptions 4.1, and within the closed, bounded, and convex set $\mathcal{C}\subset\mathbb{R}^{d}$, the gradient of $\ell$ is uniformly bounded by some constant $G\geq 0$ such that for all $z\in\mathcal{Z}$ and $\theta\in\mathcal{C}$, $\lVert\nabla\ell(z;\theta)\rVert\leq G$. We implement PSGD-R2D (Algorithms 1 and 2) with $\sigma=\frac{\Sigma\sqrt{2\log(1.25/\delta)}}{\varepsilon\delta}.$ Then for $0<\varepsilon\leq 1$ and $\delta>0$, PSGD-R2D is an $(\varepsilon,2\delta)$-unlearning algorithm for the following choices of $\Sigma$:
@@
  • If $\ell$ is convex (satisfies Assumption 4.2), then for $\eta\leq\frac{2}{L}$ we have
@@
  • If $\ell$ is $\mu$-strongly convex (satisfies Assumption 4.3), then for $\eta\leq\frac{\mu}{L^{2}}$ and $\gamma=\sqrt{1-\eta\mu}$ we have
@@
Theorem 4.5 highlights the privacy-utility-complexity tradeoff of PSGD-R2D, through the relationship between $\Sigma$, $\varepsilon$ and $K$. We observe that for all settings, $\Sigma$ decreases to zero as $K$ increases from zero to $T$, and it decays exponentially for strongly convex functions and linearly for convex functions. For general nonconvex functions, $\Sigma$ still decreases to zero as $K$ goes to $T$, but at a slower rate. This reflects the contractive, semi-contractive, and expansive properties of gradient algorithms on strongly convex, convex, and nonconvex functions respectively.
@@
If $\ell$ is $\mu$-strongly convex (satisfies Assumption 4.3), then $\Sigma$ is uniformly upper bounded as follows.
@@
Proof. See Appendix A.5.1.
We note that $\frac{n\mu\Sigma}{2\eta G}+\gamma^{T}\leq 1$, so the equation of $K$ is overall positive. For small $T$, $K$ increases at an approximately linear rate, but as $T$ grows to infinity, $K$ converges to a constant value. This implies that for a given level of of privacy $(\varepsilon,\delta)$ and noise, the number of unlearning iterations required is uniformly upper bounded even if the number of training iterations $T$ is arbitrarily large.
@@ 4.2 SGD-R2D: Rewinding with SGD
Now we consider rewinding without projection. A key challenge of analyzing SGD unlearning algorithms on an unbounded domain is that the noise and unlearning bias may also be unbounded. We therefore require a reasonable assumption that the second moment of the stochastic gradient is relatively bounded, as in Assumption 4.7.
Assumption 4.7.
For all datasets $\tilde{\mathcal{D}}\sim\mathcal{Z}$, let $g_{\mathcal{B}}(\theta)$ represent the gradient estimator constructed from a batch $\mathcal{B}\sim\tilde{\mathcal{D}}$ uniformly sampled with replacement. Then the second moment of the gradient estimator is relatively bounded with constants $B,C\geq 0$ such that for all $\theta\in\mathbb{R}^{d}$,
$\mathbb{E}[\lVert g_{\mathcal{B}}(\theta)\rVert^{2}]\leq B\lVert\nabla\mathcal{L}_{\tilde{\mathcal{D}}}(\theta)\rVert^{2}+C.$ (7)
For finite-sum problems (1), Assumption 4.7 is satisfied by strongly convex loss functions and functions that satisfy the Polyak–Łojasiewicz (PL) inequality, as established in (Khaled and Richtárik, 2020; Garrigos and Gower, 2024; Gower et al., 2021). In particular, for batch sampling with replacement, the constant $B$ can be explicitly computed from $L$ and the batch size $b$, and the constant $C$ can be estimated from how close the model is to interpolation, where perfect interpolation (or overparametrization) implies $C=0$.
To handle the unlearning bias, we observe that it can be bounded by the $\chi^{2}$-divergence between the empirical distributions of $\mathcal{D}$ and $\mathcal{D}^{\prime}$, denoted as $P_{\mathcal{D}}$ and $P_{\mathcal{D}^{\prime}}$ respectively (Lemma A.15). This conveniently allows us to exploit existing conditions on the noise like Assumption 4.7.
@@
Assumption 4.8 is satisfied if the underlying data distribution $\mathcal{Z}$ is closed and bounded.
@@
Theorem 4.9.
Suppose that the loss function $\ell$ satisfies Assumptions 4.1, 4.7, and 4.8 and we implement SGD-R2D (Algorithms 1 and 2) with $\mathcal{C}=\mathbb{R}^{d}$ and $\sigma=\frac{\Sigma\sqrt{2\log(1.25/\delta)}}{\varepsilon\delta}.$ Then for $0<\varepsilon\leq 1$ and $\delta>0$, SGD-R2D is an $(\varepsilon,2\delta)$-unlearning algorithm for the following $\Sigma$:
@@
    $\Sigma=O(((1+\eta L)^{T}-(1+\eta L)^{K})(T-K)^{1/2}).$
  • If $\ell$ is convex (satisfies Assumption 4.2), then for $\eta\leq\frac{2}{L}$ we have
    $\Sigma=O((T-K)^{3/2}).$
  • If $\ell$ is $\mu$-strongly convex (satisfies Assumption 4.3), then for $\eta\leq\frac{\mu}{L^{2}}$ and $\gamma=\sqrt{1-\eta\mu}$ we have
    $\Sigma=O((\gamma^{K}-\gamma^{T})(T-K)^{1/2}).$
For all statements, the $O(\cdot)$ notation hides dependencies on $B$, $C$, $\eta$, $L$, $\ell_{\theta_{0}}$, $m$, $n$, and $\mu$.
Proof. See Appendix A.6.